OpenStudy (anonymous):

if x= a tan^2 theta when x= 0, then theta=? and when x= a, then theta=?

hartnn (hartnn):

when x= 0 \(0= a \tan^2 \theta \\ 0=\tan \theta\) for what value of theta is tan theta =0 ???

OpenStudy (anonymous):

it means when x= 0, then theta=0

hartnn (hartnn):

correct :) when x=a \(a= a \tan^2 \theta \\ 1=\tan \theta\) for what value of theta is tan theta =1 ???

OpenStudy (anonymous):

pi/4

hartnn (hartnn):

yes, pi/4 is correct. is this problem related to integration ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

oh, then ok. otherwise, we have multiple values of theta when tan theta =0 or 1, but for integration purpose, what we did is sufficient.